OpenStudy (mathisfun13):

For the geometric sequence with a1=4 and r=5 find a5

OpenStudy (mathisfun13):

that is the geometric sequence

OpenStudy (mathisfun13):

i dont know it

OpenStudy (texaschic101):

an = a1 * r ^ n-1 r = 5 a1 = 5 n = 5 a5 = 5 * 5^(5 - 1) a5 = 5 * 5^4 a5 = 5 * 625 a5 = 3125 Fifth term is 3125

OpenStudy (texaschic101):

wait....I messed up

OpenStudy (texaschic101):

let me do it again....

OpenStudy (mathisfun13):

okay. i got 31250

OpenStudy (texaschic101):

an = a1 * r ^ n - 1 a1 = 4 r = 5 n = 5 a5 = 4 * 5^(5 - 1) a5 = 4 * 5^4 a5 = 4 * 625 a5 = 2500 How did you get 31250 ?

OpenStudy (mathisfun13):

i found it in my notes from when i did this problem before

OpenStudy (mathisfun13):

prob wrong

OpenStudy (texaschic101):

did you get n = 5 ? It has to be 5, for the 5th term. A1 = 4, r = 5. I can't see anything I did wrong.You weren't looking for the sum were you ?

OpenStudy (texaschic101):

let me check it this way... a1 = 4 a2 = 4 x 5 = 20 a3 = 20 x 5 = 100 a4 = 100 x 5 = 500 a5 = 500 x 5 = 2500 correct........it is 2500

OpenStudy (mathisfun13):

no not the sum

OpenStudy (texaschic101):

a5 = 2500